OpenStudy (anonymous):

Solve 8 = 2x + 4. x = −4 x = −1 x = 0 x = 7

OpenStudy (anonymous):

Wait the actual equation is

OpenStudy (unklerhaukus):

first: take four from both sides, then: divide both sides by two

OpenStudy (anonymous):

I'm trying to post the actual equation but the buttons for the equation draw and attach file are messing up. The actual equation is 8=2^x+4 the x+4 is an exponent

OpenStudy (unklerhaukus):

(unfortunately OS isn't working properly)

OpenStudy (anonymous):

Well you know what I mean right?

OpenStudy (unklerhaukus):

8=2^{x+4} can you rewrite the lefthand side (eight) as a power of two?

OpenStudy (anonymous):

What do you mean? :/

OpenStudy (unklerhaukus):

well if it was 9 you could rewrite it as 3^2

OpenStudy (unklerhaukus):

(that is nine as a power of three) can you rewrite eight as a power of two?

OpenStudy (anonymous):

2^2?

OpenStudy (anonymous):

4^2 sorry openstudy is running so weirdly for me

OpenStudy (unklerhaukus):

nope 2^2=4 and 4^2 = 16 2^? = 8

OpenStudy (anonymous):

3

OpenStudy (unklerhaukus):

yeah 2*2*2=8 so 8=2^3

OpenStudy (unklerhaukus):

now you can write 8=2^3=2^{x+4} comparing the indexes 3=x+4

OpenStudy (anonymous):

Thank you so much~

OpenStudy (unklerhaukus):

can you solve for x now?